OpenStudy (anonymous):

If a circle has a circumference of C, what is the area of the circle in terms of C? A=πC A=C24π A=2Cπ A=2C2

OpenStudy (danjs):

The area is A=pi*r^2, and the Circumference is C=2*pi*r the ansers have pi in them, so solve both Area and Circumference for the radius r

OpenStudy (danjs):

then you can set them equal to each other since they will both equal r

OpenStudy (danjs):

\[r = \sqrt{\frac{ A }{ \pi }}\] and \[r = \frac{ C }{ 2\pi }\]

OpenStudy (danjs):

Set the two right sides equal and solve for A

OpenStudy (danjs):

k

OpenStudy (radar):

Is their a correct choice? I think we need\[A=(C ^{2})/(4\pi)\]Am I missing something??

OpenStudy (radar):

Probably the 2nd option has a typo.